2145000 + x/858000 + x = 2
2=(2,145,000+X)/(858,000+X)
Cross multiply to get:
2(858,000 + X) = 2,145,000 + X
1,716,000 + 2X = 2,145,000 + X
Subtract X from both sides
2X- X = 2,145,000 – 1,716,000
Subtract 1,716,000 from both sides
X = 429,000

Answer by Sillygirl39452(1)

2=(2,145,000+X)/(858,000+X)
Cross multiply to get:
2(858,000 + X) = 2,145,000 + X
1,716,000 + 2X = 2,145,000 + X
Subtract X from both sides
2X- X = 2,145,000 – 1,716,000
Subtract 1,716,000 from both sides
X = 429,000